OpenStudy (anonymous):

Estimate the radius of the cesium ion, Cs+. The lattice energy of cesium chloride, CsCl, is 633 kJ/mol. For CsCl, the Madelung constant M is 1.763, and the Born exponent n is 10.7. The radius of Cl− is known to be 1.81 A˚. Express your answer in Angstroms. Can you show all working, I'm not sure where to begin? Thanks.

OpenStudy (turingtest):

the lattice energy is given by\[E_{lattice}=\frac{N_AMQ_1Q_2e^2}{4\pi\epsilon_0r_0}\left(1-\frac1n\right)\]where \(N_A\) is Avagadro's number, \(M\) is the Madelung constant, \(Q1,Q_2\) are the magnitude of the charge of the ions, and \(r_0\) is the total separation of the ion centers (which is the sum of the radii of each ion) Use the given info to solve for the equilibrium separation \(r_0\) and use the fact that the ionic radius of Cl is given to find the ionic radius of Cs.

OpenStudy (turingtest):

that energy is in kJ/mol btw, so you will have to put it back into Joules to get the radius

OpenStudy (anonymous):

oh okay thank you

OpenStudy (turingtest):

welcome!

OpenStudy (anonymous):

what do I use in the denominator?? it's 4 * 3.142 * ? * ?

OpenStudy (turingtest):

\(\epsilon_0\approx8.85\times10^{-12}\) is the permittivity of free space, and since you are solving for \(r_0\) you should have \(E_{lattice}=633\times10^3\) as the other number

OpenStudy (anonymous):

I got the answer to be 1.67!!! Thank-you for the help!!! :) I appreciate it!

OpenStudy (turingtest):

welcome!

OpenStudy (anonymous):

What is \[Q _{1} and Q _{2}\]?